A refrigerator requires 5 hp to provide 24000 kJ/h
Answer #1

24000 kJ/h = 24000*1000 / 60*60 = 6666.67 W

5 hp = 5*746 = 3730 W

Coefficent of Perfoemance = Heat rejected / work done

= 6666.67 / (3730)

= 1.787

so option c) 1.8 is the correct answer